# On uniqueness of radial potentials for given Dirichlet spectra with distinct angular momenta

### v1

@@ Abstract
We consider an inverse spectral problem for radial Schrödinger operators with singular potentials. First, we show that the knowledge of the Dirichlet spectra for infinitely many angular momenta $\ell$ satisfying a Müntz-type condition uniquely determines the potential. Then, in a neighborhood of the zero potential, we prove that the potential is uniquely determined by two Dirichlet spectra associated with distinct angular momenta in the cases $(\ell_{1},\ell_{2})=(0,1)$ and $(0,2)$. Our approach relies on an explicit analysis of the corresponding singular differential equation, combined with the classical Kneser–Sommerfeld formula. These results confirm, in the linearized setting and in these configurations, a conjecture originally formulated by Rundell and Sacks (2001).
@@ 1 Introduction
Before addressing this question, let us recall some spectral properties of the underlying operator. For each fixed angular momentum $\ell\in\mathbb{N}$ and potential $q\in L^{2}(0,1)$, the Dirichlet spectrum is given by the zeros of an entire function (the so-called regular solution; see Appendix B), and consists of a countable sequence of simple, real eigenvalues $\{\lambda_{\ell,n}(q)\}_{n\geq 1}$ tending to infinity. These properties are well established in the literature [4, 5] and form the basis for both direct and inverse spectral analysis in the radial setting.
@@
Before turning to the more delicate case of two angular momenta, let us first consider the situation where Dirichlet spectra are known for infinitely many values of $\ell$. Our first theorem is a global uniqueness result and concerns the case where one knows the Dirichlet spectra for infinitely many angular momenta $\ell$ satisfying a Müntz-type condition. In fact, this result is a direct consequence of Theorem 2.1 established in [20] in the framework of scattering theory; see also ([6], Theorem 1.4) for local uniqueness results and [8] for an extension to magnetic fields. The proof of Theorem 1.1 is given in Appendix B, where we establish a correspondence between Dirichlet spectral data and scattering phases.
@@
In the unperturbed case $q\equiv 0$, the eigenfunctions of (1.1) can be expressed explicitly in terms of Bessel functions of the first kind. Specifically, for each $\ell\in\mathbb{N}$, the eigenfunctions take the form (up to normalization) [^1] For simplicity of exposition, we shall omit this inessential normalization throughout the present paper.
@@
is dense in $L^{2}(0,1)$. Establishing this completeness property in the linearized setting is a first step in order to prove the local uniqueness of the potential near $q=0$.
Even in the more favorable case where the angular momentum $\ell$ takes infinitely many values, it is not known whether the vector space spanned by the functions $\psi_{\ell,n}^{2}(x)$ is dense in $L^{2}(0,1)$, despite the fact that the problem is expected to be highly overdetermined. Although global uniqueness can now be derived by other methods, see Appendix B, the following result remains of independent interest.
@@
Independently of the completeness result, the asymptotic behavior of the eigenvalues imposes further constraints on admissible perturbations. Specifically, as shown in [24, Proposition 3.1] (see also [23, Eq. (2.10)]), for each fixed angular momentum $\ell$, the following expansion holds uniformly on bounded subsets of $L^{2}(0,1)$:
@@
We conclude this section with the main result of the paper. It shows that, in two specific cases, the Dirichlet spectra corresponding to two distinct angular momenta uniquely determine the potential in a neighborhood of the zero potential in the $L^{2}(0,1)$ topology.
@@
For $(\ell_{1},\ell_{2})=(0,1)$ or $(0,2)$, the knowledge of the Dirichlet spectra corresponding to angular momenta $\ell_{1}$ and $\ell_{2}$ uniquely determines the potential $q\in L^{2}(0,1)$ in a $L^{2}$-neighborhood of the zero potential.
@@ 2.1 The spectral map
The following proposition, proved in [24, Theorems 2.3 and 4.1] and based on (1.2), states that the spectral map is real-analytic and provides an explicit expression for its Fréchet differential at the zero potential. The subtraction of 1 in the integrals removes the constant part, already represented by the first coordinate of $\mathcal{S}_{\ell_{1},\ell_{2}}$.
@@
Let $\ell_{1}\neq\ell_{2}$ be two fixed non-negative integers, and set $\nu_{k}=\ell_{k}+\tfrac{1}{2}$ for $k=1,2$. The spectral map
$\mathcal{S}_{\ell_{1},\ell_{2}}:L^{2}(0,1)\to\mathbb{R}\times\ell_{\mathbb{R}}^{2}(\mathbb{N})\times\ell_{\mathbb{R}}^{2}(\mathbb{N}),$
defined by
$\mathcal{S}_{\ell_{1},\ell_{2}}(q)=\left(\int_{0}^{1}q(r)\,dr,\;\big(\widetilde{\lambda}_{\ell_{1},n}(q)\big)_{n\geq 1},\;\big(\widetilde{\lambda}_{\ell_{2},n}(q)\big)_{n\geq 1}\right),$
@@
$d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}(\zeta)=\left(\int_{0}^{1}\zeta(r)\,dr,\quad\left(\int_{0}^{1}\zeta(r)\,\left(\psi_{\ell_{1},n}^{2}(r)-1\right)\,dr\right)_{n\geq 1},\quad\left(\int_{0}^{1}\zeta(r)\,\left(\psi_{\ell_{2},n}^{2}(r)-1\right)\,dr\right)_{n\geq 1}\right),$
where the eigenfunctions $\psi_{\ell,n}(r)$ are given, up to a normalization [^2] As mentioned earlier, we shall omit this inessential normalization., by
$\psi_{\ell,n}(r):=\sqrt{\frac{\pi r}{2}}\,J_{\nu}(j_{\nu,n}r),\quad\nu=\ell+\tfrac{1}{2},\quad n\geq 1.$
@@
is injective, provided that $(\ell_{1},\ell_{2})=(0,1)$ or $(\ell_{1},\ell_{2})=(0,2)$.
The proof of Theorem 2.2 is given in Sections 7–8.
As a byproduct of our analysis we obtain the following theorem in the case $\ell_{1}=0$ and arbitrary $\ell_{2}$.
@@
Let $\ell_{2}\geq 1$. Consider
$\mathcal{S}_{0,\ell_{2}}:L^{2}(0,1)\longrightarrow\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N}).$
Then the Fréchet differential of $\mathcal{S}_{0,\ell_{2}}$ at the zero potential $q=0$ has a finite-dimensional kernel.
@@
Our approach shows that any element of this kernel satisfies a homogeneous linear ODE on $(0,1)$, whose order and coefficients depend only on $\ell_{2}$. The cases $(\ell_{1},\ell_{2})=(0,1)$ and $(0,2)$ are treated in detail in this paper, and the extension to arbitrary $\ell_{2}$ follows without additional difficulty. Consequently, the kernel is spanned by finitely many fundamental solutions and is therefore finite-dimensional. ∎
This recovers the case $(\ell_{1},\ell_{2})=(0,1)$ proved in [25], and is consistent with [4] when $\ell_{1}-\ell_{2}$ is odd. We conjecture that the same conclusion holds for all pairs of distinct nonnegative integers $\ell_{1}\neq\ell_{2}$.
2.2 Proof of Theorem 1.3
In this subsection, we show how Theorem 2.2 implies Theorem 1.3. First, let us decompose $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}$ in two parts
$d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}=A_{\ell_{1},\ell_{2}}+K_{\ell_{1},\ell_{2}}$
where, denoting for any integer $\ell$, $\Phi_{\ell}(x):=\frac{\pi}{2}xJ_{\ell+\frac{1}{2}}(x)^{2}$,
$A_{\ell_{1},\ell_{2}}(\zeta)=\Bigg(\int_{0}^{1}\zeta(r)\,dr,$ $\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{1}}\!\left(\bigl(n+\tfrac{\ell_{1}}{2}\bigr)\pi r\right)-1\Bigr)\,dr\right)_{n\geq 1},$
$\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{2}}\!\left(\bigl(n+\tfrac{\ell_{2}}{2}\bigr)\pi r\right)-1\Bigr)\,dr\right)_{n\geq 1}\Bigg).$ (2.2)
Then, we have:
$K_{\ell_{1},\ell_{2}}(\zeta)=\Bigg($  $\,0,\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{1}}(j_{\ell_{1}+\frac{1}{2},n}\pi r)-\Phi_{\ell_{1}}\!\left(\bigl(n+\tfrac{\ell_{1}}{2}\bigr)\pi r\right)\Bigr)\,dr\right)_{n\geq 1},$
$\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{2}}(j_{\ell_{2}+\frac{1}{2},n}\pi r)-\Phi_{\ell_{2}}\!\left(\bigl(n+\tfrac{\ell_{2}}{2}\bigr)\pi r\right)\Bigr)\,dr\right)_{n\geq 1}\Bigg).$ (2.3)
Since $|\Phi^{\prime}_{\ell}(x)|\leq C_{\ell}$ uniformly on $\mathbb{R}$ and $\sum_{n}|j_{\ell+\frac{1}{2},n}-(n+\frac{\ell}{2})|^{2}<\infty$ for any integer value of $\ell$, we get
$\sum_{n}\|\Phi_{\ell}(j_{\ell,n}\pi\cdot)-\Phi_{\ell}((n+\frac{\ell}{2})\pi\cdot)\|_{L^{2}(0,1)}^{2}<\infty$
and thus, by Cauchy–Schwarz inequality, $K_{\ell_{1},\ell_{2}}$ is a Hilbert–Schmidt (hence compact) operator from $L^{2}(0,1)$ into $\mathbb{R}\times\ell^{2}(\mathbb{N})\times\ell^{2}(\mathbb{N})$. Now from ([23], Corollary 5.2) and Theorem A.2 in the appendix of this paper), we learn that, for $(\ell_{1},\ell_{2})=(0,1)$ or $(\ell_{1},\ell_{2})=(0,2)$,
$\left(1,\ \left(\Phi_{\ell_{1}}((n+\frac{\ell_{1}}{2})\pi\cdot)\right),\ \left(\Phi_{\ell_{2}}((n+\frac{\ell_{2}}{2})\pi\cdot)\right)\right)$
is a basis of $L^{2}(0,1)$ which in turn implies that $A_{\ell_{1},\ell_{2}}$ is an isomorphism from $L^{2}(0,1)$ onto $\mathbb{R}\times\ell^{2}(\mathbb{N})\times\ell^{2}(\mathbb{N})$. Therefore by Theorem 2.2 and the Fredholm alternative, $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}$ is also an isomorphism from $L^{2}(0,1)$ onto $\mathbb{R}\times\ell^{2}(\mathbb{N})\times\ell^{2}(\mathbb{N})$. Now, applying the inverse function theorem, we obtain Theorem 1.3.
@@ 4 Transformation Operators
$g(x)=f(x)-4\ell x^{-2\ell}\int_{0}^{x}s^{2\ell-1}f(s)\,ds$ (4.7)
and differentiate once for getting
@@
$(2\ell+1)g^{\prime}(x)+xg^{\prime\prime}(x)=(-2\ell+1)f^{\prime}(x)+xf^{\prime\prime}(x)$
and iterating $k$ times we get
@@ 5 Green’s identity via transformation operators
In this section, we shall employ the transformation operators $T_{\ell}$ to reformulate relation (3.3) in a more convenient form. First, it is straightforward to verify that,
@@
We will now employ the half-integer Bessel functions and their associated polynomials $P_{\ell}$ and $Q_{\ell}$ (see [1, 10.1.19–20]). For $\ell=0,1,2,\dots$ and $z\in\mathbb{C}$ ,
@@ 6.3 The case $\ell\geq 2$
7 Uniqueness result in the case $\ell=0$ and $\ell=1$
@@
In this section, we assume that $\zeta$ satisfies the hypotheses of Propostion 5.1 for $\ell=0$ and $\ell=1$ . For $\ell=0$ , this simply implies that $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$. For $\ell=1$ , setting $g:=T_{1}[\zeta]$, the result means that the function $g^{\prime}(x)+2g(x)$ is odd in the sense of distributions.
The goal is to reduce the problem to the study of an ordinary differential equation involving $\zeta$ only. According to Lemma 4.1 (vi), applied with $\ell=1$, and since $g=S_{1}[\zeta]$, we have:
$g^{\prime\prime}(x)=\zeta^{\prime\prime}(x)+\frac{4}{x}\,\zeta^{\prime}(x)\,.$ (7.1)
Differentiating both sides gives:
$g^{(3)}(x)=\zeta^{(3)}(x)+\frac{4}{x}\zeta^{\prime\prime}(x)-\frac{4}{x^{2}}\zeta^{\prime}(x)\,.$ (7.2)
Since $g^{\prime}(x)+2g(x)$ is odd, it follows that $g^{(3)}(x)+2g^{\prime\prime}(x)$ is also odd in the sense of distributions. Using equations (7.1) and (7.2), we deduce that the following expression is odd:
$f(x):=-\zeta^{(3)}(x)+\left(2-\frac{4}{x}\right)\zeta^{\prime\prime}(x)+\left(\frac{4}{x^{2}}+\frac{8}{x}\right)\zeta^{\prime}(x)\,.$
Letting $y:=\zeta^{\prime}(x)$, we rewrite this as
$f(x)=-y^{\prime\prime}(x)+\left(2-\frac{4}{x}\right)y^{\prime}(x)+\left(\frac{4}{x^{2}}+\frac{8}{x}\right)y(x)\,$, (7.3)
which is odd with respect to $x=\tfrac{1}{2}$, that is satisfies, f(x)=f(1x)f(x)=-f(1-x) .
Taking into account the parity of $y$, and after elementary calculations, we arrive at the following lemma:
Lemma 7.1.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Lemma 5.1 for $\ell=0$ and $\ell=1$, and set $y(x):=\zeta^{\prime}(x)$. Then $y$ satisfies the differential equation
$y^{\prime\prime}(x)+\left(\frac{2}{x}-\frac{2}{1-x}\right)y^{\prime}(x)-\left(\frac{2}{x^{2}}+\frac{2}{(1-x)^{2}}+\frac{4}{x}+\frac{4}{1-x}\right)y(x)=0$ (7.4)
in the sense of distributions on $(0,1)$.
Now, we seek symmetric solutions of the differential equation (7.4). The equation is weakly singular at both endpoints $x=0$ and $x=1$ ([28], p. 236, or [29]), and we begin by analyzing the behavior near $x=0$. The indicial equation associated with the singularity at zero is
$\rho(\rho-1)+2\rho-2=0,$
which has roots $\rho=1$ and $\rho=-2$. There is a basis of solutions near zero with one asymptotic to $x^{-2}$ and the other asymptotic to $x$. The solution which is asymptotic to $x^{-2}$ is to exclude since this would lead to a $\zeta$ which is not in $L^{2}(0,1)$. Consequently, our solution is proportional to the second one and satisfies $y(0)=0$.
A similar analysis near $x=1$ leads to the same conclusion: the admissible solution must satisfy $y(1)=0$. Hence, we restrict ourselves to functions $y\in H_{0}^{1}(0,1)$ satisfying homogeneous Dirichlet boundary conditions.
Since $f(x)$ given in (7.3) is odd and $y(x)$ is even, we have
[…]
A straightforward integration by parts, using the boundary conditions $y(0)=y(1)=0$, yields
$0=\int_{0}^{1}f(x)\,y(x)\,dx\;=\;\int_{0}^{1}\bigl(y^{\prime}(x)\bigr)^{2}\,dx\;+\;\int_{0}^{1}\left(\frac{2}{x^{2}}+\frac{8}{x}\right)\,y^{2}(x)\,dx.$ (7.5)
It follows that $y(x)=0$ almost everywhere. We conclude that $\zeta$ must be constant. But since $\zeta$ is odd with respect to $x=\tfrac{1}{2}$, we get $\zeta=0$.
We conclude this section with the following completeness result:

### v3

@@ Abstract
We consider an inverse spectral problem for radial Schrödinger operators with singular potentials. First, we show that the knowledge of the Dirichlet spectra for infinitely many angular momenta $\ell$ satisfying a Müntz-type condition uniquely determines the potential. Next, in a neighborhood of the zero potential, we prove local uniqueness from two Dirichlet spectra associated with distinct angular momenta in the cases $(\ell_{1},\ell_{2})=(0,1)\,,\ (1,2)$ and $(0,3)$ . Our approach relies on an explicit analysis of the associated singular differential equation, combined with the classical Kneser–Sommerfeld formula. These results sharpen a theorem of Carlson-Shubin (1994) and confirm, in the linearized setting and for these configurations, a conjecture originally formulated by Rundell and Sacks (2001).
@@ 1 Introduction
Before addressing this question, let us recall some spectral properties of the underlying operator. For each fixed angular momentum $\ell\in\mathbb{N}$ and potential $q\in L^{2}(0,1)$, the Dirichlet spectrum is given by the zeros of an entire function (the so-called regular solution; see Appendix A), and consists of a countable sequence of simple, real eigenvalues $\{\lambda_{\ell,n}(q)\}_{n\geq 1}$ tending to infinity. These properties are well established in the literature [5, 6] and form the basis for both direct and inverse spectral analysis in the radial setting.
@@
Before turning to the more delicate case of two angular momenta, let us first consider the situation where Dirichlet spectra are known for infinitely many values of $\ell\,$. Our first theorem is a global uniqueness result and concerns the case where one knows the Dirichlet spectra for infinitely many angular momenta $\ell$ satisfying a Müntz-type condition. In fact, this result is a direct consequence of Theorem 2.1 established in [24] in the framework of scattering theory; see also ([7], Theorem 1.4) for local uniqueness results and [9] for an extension to magnetic fields. The proof of Theorem 1.1 is given in Appendix A , where we establish a correspondence between Dirichlet spectral data and scattering phases.
@@
In the unperturbed case $q\equiv 0$, the eigenfunctions of (1.1) can be expressed explicitly in terms of Bessel functions of the first kind. Specifically, for each $\ell\in\mathbb{N}$ , the eigenfunctions take the form (up to normalization)
@@
is dense in $L^{2}(0,1)$. Establishing this completeness property in the linearized setting is a first step in order to prove the local uniqueness of the potential near q=0q=0\,.
Even in the more favorable case where the angular momentum $\ell$ takes infinitely many values, it is not known whether the vector space spanned by the functions $\psi_{\ell,n}^{2}(x)$ is dense in $L^{2}(0,1)$ , despite the fact that the problem is expected to be highly overdetermined. Although global uniqueness can now be derived by other methods, see Appendix A, the following result remains of independent interest.
@@
The proof of this theorem is given in Section 3.2.
Independently of the completeness result, the asymptotic behavior of the eigenvalues imposes further constraints on admissible perturbations. Specifically, as shown in [28, Proposition 3.1] (see also [27, Eq. (2.10)]), for each fixed angular momentum $\ell$, the following expansion holds uniformly on bounded subsets of $L^{2}(0,1)$ :
@@
We conclude this section with the main result of the paper. It shows that, in three specific cases, the Dirichlet spectra corresponding to two distinct angular momenta uniquely determine the potential in a neighborhood of the zero potential in the $L^{2}(0,1)$ topology.
@@
For $(\ell_{1},\ell_{2})\in\{(0,1),(1,2),(0,3)\}$ , the knowledge of the Dirichlet spectra corresponding to angular momenta $\ell_{1}$ and $\ell_{2}$ uniquely determines the potential $q\in L^{2}(0,1)$ in a $L^{2}$-neighborhood of the zero potential.
Remark 1.4.
The case $(\ell_{1},\ell_{2})=(0,2)$ is also investigated in this paper. We show that the differential of the spectral map (see Section 2 for its definition) is injective at $q=0$ . We conjecture that this differential is an isomorphism. If this is the case, Theorem 1.3 would also hold for $(\ell_{1},\ell_{2})=(0,2)$ by the local inversion theorem. We leave this question as an open problem.
@@ 2.1 The spectral map
The following proposition, proved in [28, Theorems 2.3 and 4.1] and based on (1.2), states that the spectral map is real-analytic and provides an explicit expression for its Fréchet differential at the zero potential.
Before stating the proposition, we introduce the following notation.
Notation 2.1.
The spherical Bessel function $j_{\ell}$ is defined by
$j_{\ell}(z)=\sqrt{\frac{\pi z}{2}}\,J_{\nu}(z)\,,\qquad\nu=\ell+\tfrac{1}{2}\,$,
where $J_{\nu}$ denotes the Bessel function of the first kind. For $n\in\mathbb{N}$ , we set
$g_{\ell,n}(x)=\frac{j_{\ell}(j_{\nu,n}x)}{\|\,j_{\ell}(j_{\nu,n}\,\cdot)\,\|_{L^{2}(0,1)}}\,.$
@@
Let $\ell_{1}\neq\ell_{2}$ be two fixed non-negative integers. The spectral map
$\mathcal{S}_{\ell_{1},\ell_{2}}:L^{2}(0,1)\to\mathbb{R}\times\ell_{\mathbb{R}}^{2}(\mathbb{N})\times\ell_{\mathbb{R}}^{2}(\mathbb{N})$
@@
$d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}(\zeta)=\left(\langle\zeta,1\rangle\,,\quad\bigl(\langle\zeta,g_{\ell_{1},n}^{2}-1\rangle\bigr)_{n\geq 1},\quad\bigl(\langle\zeta,g_{\ell_{2},n}^{2}-1\rangle\bigr)_{n\geq 1}\right)\,$,
where $\langle\cdot,\cdot\rangle$ denotes the scalar product in $L^{2}(0,1)$ .
@@
is injective, provided that $(\ell_{1},\ell_{2})\in\{(0,1)\,,\,(0,2)\,,\,(1,2)\,,\,(0,3)\}$ .
The proof of Theorem 2.2 is given in Sections 7-9.
As a byproduct of our analysis, we obtain the following theorem in the case $\ell_{1}=0$ or $1$ and arbitrary $\ell_{2}$ .
@@
Let $\ell_{1}=0$ or $1$ , and $\ell_{2}\geq 1$ . Consider
$\mathcal{S}_{\ell_{1},\ell_{2}}:L^{2}(0,1)\longrightarrow\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\,.$
Then the Fréchet differential of $\mathcal{S}_{\ell_{1},\ell_{2}}$ at the zero potential $q=0$ has a finite-dimensional kernel.
@@
Our approach shows that any element of this kernel satisfies a homogeneous linear ODE on $(0,1)$, whose order and coefficients depend only on $\ell_{2}$ . The cases $(\ell_{1},\ell_{2})=(0,1)$ , $(0,2)$ , $(1,2)$ and $(0,3)$ are treated in detail in this paper, and the extension to arbitrary $\ell_{2}$ follows without additional difficulty. Consequently, the kernel is spanned by finitely many fundamental solutions and is therefore finite-dimensional. ∎
Remark 2.4.
This result extends Theorem 1 of Shubin Christ [29] in the case $(\ell_{1},\ell_{2})=(0,1)$ , as well as the result of Carlson–Shubin [5, Theorem 1.1] in the case where $\ell_{1}-\ell_{2}$ is odd, at the level of the Fréchet differential at the zero potential $q=0$ . In these works, the authors proved that the Fréchet differential $d_{q}\mathcal{S}_{\ell_{1},\ell_{2}}$ has a finite-dimensional kernel for arbitrary potentials $q$ . We conjecture that the same conclusion holds for all pairs of distinct nonnegative integers $\ell_{1}\neq\ell_{2}$ .
We conclude this section with the following result, whose proof is given in Appendix B.
Theorem 2.5 (Closed range).
Let $(\ell_{1},\ell_{2})\in\{(0,1)\,,\,(1,2)\,,\,(0,3)\}$ . Then the differential of the spectral map at the origin, $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}$ , has closed range. More precisely, in the case $(\ell_{1},\ell_{2})=(0,1)$ , it is surjective, and its range coincides with the entire space
$\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\,.$
Consequently, in the case $(0,1)$ , the standard local inverse function theorem in Banach spaces, together with Theorem 2.2 and Theorem 2.5, implies Theorem 1.3. More precisely, in this case the spectral map provides a local real-analytic coordinate system.
The cases $(1,2)$ and $(0,3)$ are slightly different. For this reason, we recall below the following local injectivity result, which is a direct consequence of the mean value theorem and the open mapping theorem (see, for instance, [1], Theorem 2.5.10).
Proposition 2.6 (Local injectivity).
Let $X$ and $Y$ be Banach spaces, and let
$\mathcal{S}:U\subset X\longrightarrow Y$
be a $\mathcal{C}^{1}$ map defined on an open neighborhood $U$ of a point $x_{0}\in X$ . Assume that the Fréchet differential $d_{x_{0}}\mathcal{S}\,:\,X\to Y$ is injective and has closed range. Then there exists a neighborhood $V\subset U$ of $x_{0}$ such that $\mathcal{S}$ is injective on $V$.
Consequently, Theorem 1.3 follows from the previous proposition together with Theorem 2.2 and Theorem 2.5.
@@ 3.1 An integral Green identity
Remark 3.1.
The integral kernel appearing in (3.3) coincides with the gradient with respect to $q$ of the so-called regular solution evaluated at $q=0$ and $x=1$ (see [28], Proposition 2.2). This observation is consistent with Lemma A.3 of the present paper.
@@ 4 Transformation Operators
$g(x)=f(x)-4\ell x^{-2\ell}\int_{0}^{x}s^{2\ell-1}f(s)\,ds\,$, (4.7)
and differentiate once to obtain
@@
$(2\ell+1)g^{\prime}(x)+xg^{\prime\prime}(x)=(-2\ell+1)f^{\prime}(x)+xf^{\prime\prime}(x)\,$,
and iterating $k$ times we obtain
@@ 5 Green’s identity via transformation operators
In this section, we shall use the transformation operators $T_{\ell}$ to reformulate relation (3.3) in a more convenient form. First, it is straightforward to verify that,
@@
We will now employ the half-integer Bessel functions and their associated polynomials $P_{\ell}$ and $Q_{\ell}$ (see [2, 10.1.19–20]). For $\ell=0,1,2,\dots$ and $z\in\mathbb{C}$ ,
@@ 6.3 The case $\ell\geq 2$
Remark 6.8.
The polynomials $A_{\ell}$ are well known. More precisely, if $\theta_{\ell}$ denotes the reverse Bessel polynomial (also called the reverse Bessel polynomial of degree $\ell$) , then
$A_{\ell}(t)=\theta_{\ell}\!\left(-\frac{t}{2}\right),\qquad\ell\in\mathbb{N},$
see, e.g., [22, 18.34]. In particular, $A_{\ell}$ admits the explicit expansion
$A_{\ell}(t)=\sum_{k=0}^{\ell}\frac{(\ell+k)!}{(\ell-k)!\,k!}\,\frac{(-1)^{\ell-k}}{2^{\ell}}\,t^{\ell-k}\,.$ (6.3)
Moreover, $A_{\ell}$ satisfies the following second order differential equation
$t\,A_{\ell}^{\prime\prime}(t)+\bigl(t-2\ell\bigr)A_{\ell}^{\prime}(t)-\ell\,A_{\ell}(t)=0\,.$ (6.4)
7 Uniqueness result in the case $(\ell_{1},\ell_{2})=(0,1)$
@@
In this section, we assume that $\zeta$ satisfies the hypotheses of Proposition 5.1 for $\ell=0$ and $\ell=1$ and we aim to show that $\zeta$ vanishes almost everywhere on $(0,1)$ . For $\ell=0$ , this assumption simply implies that $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$ . For $\ell=1$ , it means by Theorem 6.6 that the function $A_{1}(D)\bigl[T_{1}[\zeta]\bigr]$ is odd in the sense of distributions.
The aim is to prove that $y=\zeta^{\prime}$ is a solution of a linear second-order differential equation on $(0,1)$ and satisfies the conditions
$y\!\left(\tfrac{1}{2}\right)=y^{\prime}\!\left(\tfrac{1}{2}\right)=0\,.$
It then follows that $y$ vanishes identically on $(0,1)$ , and consequently $\zeta$ vanishes almost everywhere on $(0,1)$ , since it is an odd function.
A straightforward computation yields
$2A_{1}(D)\bigl[T_{1}[\zeta]\bigr](x)=2A_{1}(D)\bigl[S_{1}[\zeta]\bigr](x)=-\zeta^{\prime}(x)+\left(2-\frac{4}{x}\right)\zeta(x)-4(2x-1)\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt\,.$ (7.1)
We further compute, for all $x\in(0,1)$ ,
$G(x):=D^{2}A_{1}(D)\bigl[T_{1}[\zeta]\bigr](x)=A_{1}(D)\bigl[D^{2}S_{1}[\zeta]\bigr](x)=A_{1}(D)\left[\zeta^{\prime\prime}+\frac{4}{x}\zeta^{\prime}\right](x)$
where we used Lemma 4.1 (iv) with $l=1$. Setting $y:=\zeta^{\prime}$ , we obtain for all $x\in(0,1)$ ,
$G(x)=-\frac{1}{2}y^{\prime\prime}(x)+\left(1-\frac{2}{x}\right)y^{\prime}(x)+\left(\frac{4}{x}+\frac{2}{x^{2}}\right)y(x)\,.$ (7.2)
Recalling that $G$ is odd on $(0,1)$ , it satisfies
G(x)+G(1x)=0G(x)+G(1-x)=0 (7.3)
in the sense of distributions. Since $y$ is even, the identities (7.2) and (7.3) imply that $y$ satisfies a linear second-order differential equation on $(0,1)$. Moreover, since $\zeta$ is odd, $y^{\prime}=\zeta^{\prime\prime}$ is also odd and $y^{\prime}\left(\frac{1}{2}\right)=0\,$. Finally, evaluating (7.1) at $x=\frac{1}{2}$, we obtain
$y\!\left(\tfrac{1}{2}\right)=\zeta^{\prime}\!\left(\tfrac{1}{2}\right)=-2A_{1}(D)\bigl[T_{1}[\zeta]\bigr]\!\left(\tfrac{1}{2}\right)-6\zeta\!\left(\tfrac{1}{2}\right)=0$
since both $\zeta$ and $A_{1}(D)\bigl[T_{1}[\zeta]\bigr]$ are odd functions on $(0,1)$.
The function $y$ thus satisfies a linear second-order differential equation on $(0,1)$ , together with the conditions $y\!\left(\tfrac{1}{2}\right)=0$ and $y^{\prime}\!\left(\tfrac{1}{2}\right)=0$ . By the Cauchy–Lipschitz theorem, we conclude that $y\equiv 0$ . Therefore $\zeta^{\prime}=y=0$, so $\zeta$ is constant. Since $\zeta$ is odd with respect to $x=\tfrac{1}{2}$ , this constant must vanish, and $\zeta\equiv 0$ . This completes the proof of the following completeness result.